1. write as a fraction in simplified form.

152% 7 3/4%

2. wanda borrowed &3000 at a simple interset rate of 16%.
a.How much interest did she pay in six months?
b.How much interest did she pay in two years?

Answer by checkley71(8403)

152/100=76/50=38/25
(7 3/4)/100=7.75/100=1.55/20=.31/4
----------------------------------------------
3000*.16*.5=240 INTEREST FOR 6 MONTHS.
3000*.16*2=960 INTEREST FOR 2 YEARS.